Coffee costing $4.80 per pound is blended with coffee costing $3.60 per
pound. How many pounds of each must be mixed to obtain 25
pounds of a blend that costs an average of $4.25 per pound?

Answer by stanbon(75887)

Coffee costing $4.80 per pound is blended with coffee costing $3.60 per
pound. How many pounds of each must be mixed to obtain 25
pounds of a blend that costs an average of $4.25 per pound?
------------------
Equations::
Quantity:: f + t = 25 lbs
Value::: 4.8f+3.6t = 4.25*25
------
Modify for elimination::
4.8f + 4.8t = 4.8^25
4.8f + 3.6t = 4.25*25
-----------------------------
Subtract and solve for "t":
1.2t = 0.55*25
t = 11.46 lbs (amt. of $3.60 coffee needed)
---
f = 25-11.46 = 13.54 lbs (amt. of $4.80 coffee needed)
--------------
Cheers,
Stan H.
---------------